OpenStudy (anonymous):

PLEASE HELP!! :( Solve the following quadratic equations. (hint: use factorization) p^2+9p-22=0 URGENT please!!

OpenStudy (anonymous):

p^2+11p-2p-22=0 p(p+11)-2(p+11)=0 (p+11)(p-2)=0 p=-11 or p=2

OpenStudy (anonymous):

This one too please, the same question but, n^2+8n+22=7

OpenStudy (anonymous):

n^2+8n+22-7=0 n^2+8n+15=0 n^2+5n+3n+15=0 n(n+5)+3(n+5)=0 (n+3)(n+5)=0 n=-3 or n=-5

OpenStudy (anonymous):

Thanks Thanks! :D This one? :D (2m+3)(4m+3)=0

OpenStudy (anonymous):

set each factor=0 2m+3=0=>n=-3/2 or 4m+3=0=>m=-3/4

OpenStudy (anonymous):

i dont get it! :S

OpenStudy (anonymous):

how did you do it?